OpenStudy (anonymous):

what is 5\6 divided by 4\6

pooja195 (pooja195):

https://www.desmos.com/calculator

pooja195 (pooja195):

Its a calculator use it :)

OpenStudy (anonymous):

thx

Nnesha (nnesha):

 is this your question ??

OpenStudy (anonymous):

yes

pooja195 (pooja195):

just subtract 5-4 and keep the denominator the same

OpenStudy (anonymous):

ok thx

Nnesha (nnesha):

okay so you have to change division to multiplication \[\frac{ \frac{ a }{ b} }{ \frac{ c }{ d } } =\frac{ a }{ b } \times \frac{ d }{ c }\]